# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ 2.3 Hyperparameter Choice
We next proceed to estimate $\tau^{2}$, which can be interpreted as the global shrinkage parameter of the loadings. For this, we employ an empirical Bayes approach. Since $\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2},\tau^{2}\sim N_{k}(0,\tau^{2}\widetilde{\sigma}_{j}^{2}\mathbb{I}_{k})$ for $j=1,\ldots,p$ a priori, we obtain an estimate of $\tau^{2}$ by conditioning on $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})$ and then maximizing the conditional likelihood, leading to $\widehat{\tau}_{0}^{2}=\sum_{j=1}^{p}\left({\|\widetilde{\lambda}_{j}\|_{2}^{2}}/{\widetilde{\sigma}_{j}^{2}}\right)/(kp).$ Following from the pseudo-posterior updates (10) and (11), we simply estimate $k$ by $\widehat{k}$ as before, $\|\widetilde{\lambda}_{j}\|_{2}^{2}$ by $\mathcal{L}_{j}^{2}=\|U^{\top}y^{(j)}\|_{2}^{2}/n$, and $\widetilde{\sigma}_{j}^{2}$ by $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$. We remark here that $U$ is dependent on the estimate $\widehat{k}$. This leads us to the plug-in estimate of $\tau^{2}$ given by
@@
The consistency of the point estimators $\mathcal{L}_{j}^{2}$ and $\mathcal{V}_{j}^{2}$ for $1\leq j\leq p$ is proven in Section E, Lemma E.4 of the Supplementary Material. Lastly, we observed that inference using the CC-FABLE procedure is not sensitive to $(\gamma_{0},\delta_{0}^{2})$, and thus we use the default values $\gamma_{0}=\delta_{0}^{2}=1$ in practice. Once the number of factors $k$ and the common variance $\tau^{2}$ have been estimated using $\widehat{k}$ and $\widehat{\tau}^{2}$, respectively, we then proceed to obtain pseudo-posterior samples of $\Psi$ using CC-FABLE as described before.
2.4 Final Algorithm
In this subsection, we provide an algorithm for implementing CC-FABLE on a given data set to obtain pseudo-posterior samples of the high-dimensional covariance matrix. We first tune the relevant hyperparameters $k$ and $\tau^{2}$ as in Section 2.3 and proceed to obtain coverage-corrected samples as in Section 2.2. The relevant steps are described in Algorithm 1.
Input: The data matrix $\mathbf{Y}\in\mathbb{R}^{n\times p}$, number of Monte Carlo (MC) samples $N_{0}$, inverse-gamma hyperparameters $(\gamma_{0},\sigma_{0}^{2}),$ and the upper bound on the cumulative singular value proportion ${S}_{0}$. Let $r=n\wedge p$.
Step 1: Compute the SVD of $\mathbf{Y}$ as $\mathbf{Y}=U^{*}D^{*}V^{*\top}$ with $U\in\mathbb{R}^{n\times r},D\in\mathbb{R}^{r\times r},V\in\mathbb{R}^{p\times r}$, and $D=\mbox{diag}(s_{1},\ldots,s_{r})$ such that $s_{1}\geq s_{2}\geq\ldots\geq s_{r}\geq 0.$ Step 2: Let $\widehat{k}={\mbox{argmin}}\,\{\mbox{JIC}(k^{*})\,|\,k^{*}=1,\ldots,\mathcal{K}_{0}\},$ where $\mathcal{K}_{0}=\mbox{argmin}_{1\leq\mathcal{K}\leq r}\left(\sum_{j=1}^{\mathcal{K}}s_{j}\right)/\left(\sum_{j=1}^{r}s_{j}\right).$ Step 3: Let $U$ consist of the first $\widehat{k}$ columns of $U^{*}$. For $1\leq j\leq p$, let $y^{(j)}$ denote the $j$th column of $\mathbf{Y}$, and obtain $\mathcal{L}_{j}^{2}=\|U^{\top}y^{(j)}\|_{2}^{2}/n$ and $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n.$ Step 4: Estimate $\tau^{2}$ by
$\begin{split}b_{uv}&=\left\{\begin{array}[]{ll}\left(1+\dfrac{\|\mu_{u}\|_{2}^{2}\|\mu_{v}\|_{2}^{2}+(\mu_{u}^{\top}\mu_{v})^{2}}{\mathcal{V}_{u}^{2}\|\mu_{v}\|^{2}+\mathcal{V}_{v}^{2}\|\mu_{u}\|_{2}^{2}}\right)^{1/2},&\text{if $u\neq v,$}\\ \left(1+\dfrac{\|\mu_{u}\|_{2}^{2}}{2\mathcal{V}_{u}^{2}}\right)^{1/2},&\text{if $u=v,$}\end{array}\right.\end{split}$
and let $\rho=\overline{b}=\left(\sum_{1\leq u\leq v\leq p}b_{uv}\right)/\{p(p+1)/2\}$ be the coverage-correction factor. Step 6: Let $\gamma_{n}=\gamma_{0}+n$. For $j=1,\ldots,p$, evaluate
Step 7: For each $t=1,\ldots,N_{0}$, independently sample $({\lambda}_{C,j}^{(t)},{\sigma}_{C,j}^{(t)2})$ for $j=1,\ldots,p$ from
${\sigma}_{C,j}^{(t)2}\sim\mbox{IG}\left(\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),\qquad{\lambda}_{C,j}^{(t)}\mid{\sigma}_{C,j}^{(t)2}\sim N_{k}\left(\mu_{j},\dfrac{\rho^{2}{\sigma}_{C,j}^{(t)2}}{n+\widehat{\tau}^{-2}}\mathbb{I}_{k}\right).$
Form ${\Lambda}_{C}^{(t)}=[{\lambda}_{C,1}^{(t)},\ldots,{\lambda}_{C,p}^{(t)}]^{\top}$ and ${\Sigma}_{C}^{(t)}=\mbox{diag}({\sigma}_{C,1}^{(t)2},\ldots,{\sigma}_{C,p}^{(t)2})$. Step 8: For each $t=1,\ldots,N_{0}$, let $L_{C}^{(t)}=\Lambda_{C}^{(t)}\Lambda_{C}^{(t)\top}$, and compute the $t$th coverage-corrected sample as
$\Psi_{C}^{(t)}=L_{C}^{(t)}+\Sigma_{C}^{(t)}.$
Output: The $N_{0}$ MC samples of the covariance matrix $\Psi_{C}^{(1)},\ldots,\Psi_{C}^{(N_{0})}$.
Algorithm 1 Steps to obtain $N_{0}$ pseudo-posterior samples of the covariance matrix using CC-FABLE.
@@ 3.1 Setup and Assumptions
In this section, we provide theoretical guarantees of the proposed CC-FABLE procedure in estimating the high-dimensional covariance matrix. Most of the existing literature (Pati et al.,, 2014; Bhattacharya and Dunson,, 2011) on obtaining provable guarantees for Bayesian factor models utilizes the machinery of Bayes’ rule in deriving posterior contraction rates. There is a clear need for further theoretical work on uncertainty quantification in such models. Xie and Wu, (2023) consider a signal-plus-noise model and derive theoretical properties of a quasi-posterior based approach, with guarantees on asymptotic frequentist coverage of the quasi-posterior intervals. However, their model only considers a square signal matrix, with the underlying focus different from high-dimensional covariance estimation. In contrast, the proposed method carries out inference on the underlying covariance matrix, without directly relying on Bayes’ rule to obtain the pseudo-posterior samples, which makes theoretical analysis of the proposed approach challenging. We overcome this challenge by leveraging on a blessing of dimensionality phenomenon which requires both $n$ and $p=p_{n}$ to grow, providing results on both the pseudo-posterior contraction and uncertainty quantification of our approach. The blessing of dimensionality phenomenon allows accurate estimation of the latent factor subspace up to rotational ambiguity, which serves as an important component behind the theoretical results. The proofs of all the theorems can be found in Sections B-D of the Supplementary Material.
@@
where $\epsilon_{i}\overset{iid}{\sim}N_{p}(0,\Sigma_{0})$ for $i=1,\ldots,n$ with $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2}),$ and $\eta_{0i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ for $i=1,\ldots,n$. Here, $\Lambda_{0}=[\lambda_{01},\ldots,\lambda_{0p}]^{\top}$ is the true factor loadings matrix while $\eta_{0i}$ are the true latent factors; integrating them out provides us $y_{i}\overset{iid}{\sim}N_{p}(0,\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0})$ for $i=1,\ldots,n$ as the marginal distribution of the data. Let $\mathbf{Y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n\times p}$ and $M_{0}=[\eta_{01},\ldots,\eta_{0n}]^{\top}\in\mathbb{R}^{n\times k}$ be the data matrix and the true matrix of latent factors, respectively, so that the true data generating model may be written as
@@
where $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}.$ Our primary goal is the inference for the covariance matrix $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}.$
For a matrix $A\in\mathbb{R}^{n_{1}\times n_{2}}$, suppose the singular values of $A$ are given by $s_{1}(A)\geq\ldots\geq s_{n_{1}\wedge n_{2}}(A)$; let $\|A\|=\underset{\|x\|_{2}=1}{\sup}\|Ax\|_{2}=s_{1}(A)$ denote the operator norm of $A$ and $\|A\|_{\infty}=\max_{ij}|A_{ij}|$ denote the max-norm of $A$, with $A_{ij}$ denoting the $ij$th entry of $A$. For two non-negative sequences $a_{m},b_{m}$, we say $a_{m}\asymp b_{m}$ if $a_{m}=\mathcal{O}(b_{m})$ and $b_{m}=\mathcal{O}(a_{m})$ as $m\to\infty$. We assume the following conditions on the true data-generating model:
@@
$s_{k}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$ as $n\to\infty$, $\|\Lambda_{0}\|_{\infty}<\infty$, and $\underset{1\leq j\leq p_{n}}{\min}\|\lambda_{0j}\|_{2}\geq c_{1}$ for some constant $c_{1}>0$.
@@
The true error variances satisfy $\underset{1\leq j\leq p_{n}}{\max}\sigma_{0j}^{2}=\mathcal{O}(1)$ and $\underset{1\leq j\leq p_{n}}{\min}\sigma_{0j}^{2}\geq c_{2}$ for some constant $c_{2}>0$.
@@
Such assumptions are standard in the existing literature on asymptotic theoretical properties of latent factor models. Assumption 3.1 allows the number of dimensions $p_{n}$ to scale as any polynomial function of $n$. Assumption 3.2 ensures that the true loadings matrix $\Lambda_{0}$ is well-conditioned, with the low-rank portion $\Lambda_{0}\Lambda_{0}^{\top}$ identifiable from noise in the asymptotic regime, and also ensures that none of the columns of $\mathbf{Y}$ consist purely of noise. Assumptions 3.3 and 3.4 assume the scalar error variances and model hyperparameters are finite. For our theoretical requirements, we will assume that $k$ and $\tau^{2}$ are known and fixed. To ease exposition, we will often suppress the dependence of $p=p_{n}$ on $n$ and express it when needed.
@@
where $U\in\mathbb{R}^{n\times k},V\in\mathbb{R}^{p\times k}$ have orthonormal columns, and $D\in\mathbb{R}^{k\times k}$ contain the $k$ largest singular values of $\mathbf{Y}$. We first provide a result showcasing the blessing of dimensionality when estimating $U_{0}U_{0}^{\top}$ by $UU^{\top}$, which forms a key part of the results that follow. Let us denote the induced CC-FABLE pseudo-posterior measure, the true data-generating measure, and the expectation under the true data-generating measure by ${\Pi}_{C}$, $P_{0}$, and $E_{0}$, respectively.
@@
Suppose Assumptions 3.13.4 hold. Then, there exists a constant $G_{1}>0$ such that
@@
The proof of Proposition 3.5 is available in Section E.1 of the Supplementary Material. First, in Theorem 3.6, we provide psuedo-posterior contraction rates when estimating $\Lambda_{0}\Lambda_{0}^{\top},\Sigma_{0},$ and $\Psi_{0}$ using $L_{C},{\Sigma}_{C}$, and ${\Psi}_{C}=L_{C}+\Sigma_{C}$, respectively, obtained from the CC-FABLE methodology in Section 2.2. Later, using Theorems 3.7 and 3.8, we justify the uncertainty quantification of the entrywise elements of $\Psi_{0}$ via pseudo-posterior credible intervals obtained from CC-FABLE.
3.2 Pseudo-posterior Contraction Rates
@@
Suppose Assumptions 3.13.4 hold. Then, as $n\to\infty$,
@@
    There exists a constant $D_{1}>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\dfrac{\|L_{C}-\Lambda_{0}\Lambda_{0}^{\top}\|}{\|\Lambda_{0}\Lambda_{0}^{\top}\|}>D_{1}\left(\dfrac{1}{\sqrt{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
@@
    There exists a constant $D_{2}>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\|\Sigma_{C}-\Sigma_{0}\|>D_{2}\left(\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}\right)\right\}\right]\to 0.$
@@
    There exists a constant $D>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\dfrac{\|\Psi_{C}-\Psi_{0}\|}{\|\Psi_{0}\|}>D\left(\dfrac{1}{\sqrt{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
Theorem 3.6 shows pseudo-posterior concentration of the relevant quantities of the CC-FABLE procedure. The relative errors when estimating each of these true parameters of interest converge to 00 at the rate $n^{-1/2}+p_{n}^{-1/2}$, which showcases the blessing of dimensionality acting on the FABLE procedure. One obtains pseudo-posterior concentration if and only if both the number of samples $n$ and the number of dimensions $p_{n}$ increase. The proof of Theorem 3.6 is available in Section B of the Supplementary Material.

### v3

@@ 2.4.2 Tuning $\tau^{2}$
After estimating $k$, we next estimate $\tau^{2}$, which can be interpreted as the global shrinkage parameter of the loadings. For this, we employ an empirical Bayes (EB) approach. Since $\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2},\tau^{2}\sim N_{k}(0,\tau^{2}\widetilde{\sigma}_{j}^{2}\mathbb{I}_{k})$ for $j=1,\ldots,p$ a priori, we obtain an estimate of $\tau^{2}$ by conditioning on $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})$ and then maximizing the conditional likelihood, leading to $\widehat{\tau}_{0}^{2}=\sum_{j=1}^{p}\left({\|\widetilde{\lambda}_{j}\|_{2}^{2}}/{\widetilde{\sigma}_{j}^{2}}\right)/(kp).$ Following from (12) and (13), we simply estimate $k$ by $\widehat{k}$ as before, $\|\widetilde{\lambda}_{j}\|_{2}^{2}$ by $\mathcal{L}_{j}^{2}=\|U^{\top}y^{(j)}\|_{2}^{2}/n$, and $\widetilde{\sigma}_{j}^{2}$ by $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$, where $U$ is dependent on $\widehat{k}$. This leads us to the plug-in estimate of $\tau^{2}$, given by
@@
The consistency of the point estimators $\mathcal{L}_{j}^{2}$ and $\mathcal{V}_{j}^{2}$ for $1\leq j\leq p$ is proven in Section E, Lemma E.4 of the Supplementary Material. As long as $\widehat{k}\to k$ as $n\to\infty$ and we assume that $(1/p)\sum_{j=1}^{p}\mathcal{L}_{j}^{2}/\mathcal{V}_{j}^{2}$ converges to a positive and finite quantity $\|\lambda_{0}\|_{2}^{2}/\sigma_{0}^{2}$ as $p\to\infty$, we obtain $\widehat{\tau}^{2}\to(1/k)\left(\|\lambda_{0}\|_{2}^{2}/\sigma_{0}^{2}\right)$. Here, $\|\lambda_{0}\|_{2}^{2}/\sigma_{0}^{2}$ may be interpreted as the true average signal-to-noise ratio (SNR). Thus, the EB estimate $\widehat{\tau}^{2}$ automatically adapts to the appropriate SNR underlying the data, with lower SNR settings producing smaller $\widehat{\tau}^{2}$ that lead to smaller loadings estimates. In simulations, we obtained better estimation accuracy using the EB estimate $\widehat{\tau}^{2}$ instead of assuming a vague prior for the loadings $\widetilde{\lambda}_{j}$ obtained by letting $\tau^{2}\to\infty$.
Lastly, we observed that the FABLE procedure is not sensitive to the hyperparameters of the prior for the variances as in (9), namely $(\gamma_{0},\delta_{0}^{2})$. Thus, we use the default values $\gamma_{0}=\delta_{0}^{2}=1$ in practice. Once the number of factors $k$ and the common variance $\tau^{2}$ have been estimated using $\widehat{k}$ and $\widehat{\tau}^{2}$, respectively, we then proceed to obtain FABLE-posterior samples of $\Psi$ as described earlier. We provide a summary of implementing the proposed method in Algorithm 1, available in Section I of the Supplementary Material.
@@ 3.1 Setup and Assumptions
In this Section, we provide theoretical guarantees on the FABLE-posterior for high-dimensional covariance matrices. Most of the existing literature (Pati et al.,, 2014; Bhattacharya and Dunson,, 2011; Ročková and George,, 2016; Srivastava et al.,, 2017) on frequentist asymptotic guarantees for Bayesian factor models focuses on posterior contraction rates. Theorems on frequentist coverage of Bayesian credible sets for such models are lacking. Xie and Wu, (2024) consider a signal-plus-noise model and derive theoretical properties of a quasi-posterior based approach, with guarantees on asymptotic frequentist coverage of the quasi-posterior intervals. However, their model only considers a square signal matrix, with the underlying focus different from high-dimensional covariance estimation. Due to the (innovative) form of the FABLE-posterior, established tools for showing Bernstein-von Mises-type results in Bayesian models cannot be used. We overcome this challenge by leveraging on a blessing of dimensionality phenomenon which requires both $n$ and $p=p_{n}$ to grow, providing results on both contraction and uncertainty quantification of the FABLE-posterior. The blessing of dimensionality is key for accurate estimation of the latent factor subspace up to rotational ambiguity. The proofs of all the theorems in this Section can be found in Sections B-D of the Supplementary Material.
@@
where $\epsilon_{i}\overset{iid}{\sim}N_{p}(0,\Sigma_{0})$ for $i=1,\ldots,n$ with $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2}),$ and $\eta_{0i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ for $i=1,\ldots,n$. Here, $\Lambda_{0}=[\lambda_{01},\ldots,\lambda_{0p}]^{\top}$ is the true factor loadings matrix while $\eta_{0i}$ are the true latent factors; integrating them out provides $y_{i}\overset{iid}{\sim}N_{p}(0,\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0})$ for $i=1,\ldots,n$ as the marginal distribution of the data. Let $\mathbf{Y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n\times p}$ and $M_{0}=[\eta_{01},\ldots,\eta_{0n}]^{\top}\in\mathbb{R}^{n\times k}$ be the data matrix and the true matrix of latent factors, respectively, so that the true data generating model may be written as
@@
where $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}.$ Our primary goal is inference for the covariance matrix $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}.$
For a matrix $A\in\mathbb{R}^{n_{1}\times n_{2}}$, we denote its singular values by $s_{1}(A)\geq\ldots\geq s_{n_{1}\wedge n_{2}}(A)$. Let $\|A\|=\underset{\|x\|_{2}=1}{\sup}\|Ax\|_{2}=s_{1}(A)$ denote the operator norm of $A$ and $\|A\|_{\infty}=\max_{ij}|A_{ij}|$ denote the max-norm of $A$, with $A_{ij}$ denoting the $(i,j)$th entry of $A$. For two sequences $a_{m},b_{m}\geq 0$, we say $a_{m}\asymp b_{m}$ if $a_{m}=\mathcal{O}(b_{m})$ and $b_{m}=\mathcal{O}(a_{m})$ as $m\to\infty$. We assume the following conditions on the true data-generating model:
@@
$s_{k}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$ as $n\to\infty$, $\|\Lambda_{0}\|_{\infty}=O(1)$, and $\underset{1\leq j\leq p_{n}}{\min}\|\lambda_{0j}\|_{2}\geq c_{1}$ for some constant $c_{1}>0$.
@@
The true error variances satisfy $\underset{1\leq j\leq p_{n}}{\max}\sigma_{0j}^{2}=\mathcal{O}(1)$ and $\underset{1\leq j\leq p_{n}}{\min}\sigma_{0j}^{2}\geq c_{2}$ for some finite constant $c_{2}>0$.
@@
Assumption 3.5.
We have $\sqrt{n}/p_{n}\to 0$ as $n\to\infty$.
Such assumptions are standard in the literature on the asymptotic properties of latent factor models (Pati et al.,, 2014; Bhattacharya and Dunson,, 2011; Xie et al.,, 2022; Ročková and George,, 2016). Assumption 3.1 allows the number of dimensions $p_{n}$ to scale as any polynomial function of $n$. Assumption 3.2 ensures that the true loadings matrix $\Lambda_{0}$ is well-conditioned, with the low-rank portion $\Lambda_{0}\Lambda_{0}^{\top}$ identifiable from noise in the asymptotic regime, and also ensures that none of the columns of $\mathbf{Y}$ consist purely of noise. Assumptions 3.3 and 3.4 assume the scalar error variances and model hyperparameters are finite. Assumption 3.5 imposes a lower bound on the rate at which $p_{n}$ increases and is necessary for the “blessing of dimensionality” to take effect rapidly enough so that asymptotically accurate UQ can be obtained. For our theoretical requirements, we will assume that $k$ and $\tau^{2}$ are known and fixed. To ease exposition, we will often suppress the dependence of $p=p_{n}$ on $n$ and express it when needed.
Although Assumption 3.1 is sufficient for posterior concentration of relative estimation errors with FABLE, we require the stronger Assumption 3.5 to obtain accurate UQ guarantees for entrywise FABLE credible intervals. Compared to assumptions in the literature on matrix denoising theory, such as $n/p_{n}\to c\in(0,1)\text{ or }(0,\infty)$ (Ledoit and Wolf,, 2004; Wang and Fan,, 2017; Hong et al.,, 2023; Benaych-Georges and Nadakuditi,, 2012), Assumption 3.5 is substantially weaker. It is common to assume stronger conditions to obtain UQ guarantees than those needed for consistency or rate results for high-dimensional covariance or precision estimation (Van de Geer et al.,, 2014; Jankova and Van de Geer,, 2015; Ning and Liu,, 2017; Yao and Lopes,, 2023).
@@
where $U\in\mathbb{R}^{n\times k},V\in\mathbb{R}^{p\times k}$ have orthonormal columns, and $D\in\mathbb{R}^{k\times k}$ contain the $k$ largest singular values of $\mathbf{Y}$. We first provide a result showcasing the blessing of dimensionality when estimating $U_{0}U_{0}^{\top}$ by $UU^{\top}$, which forms a key part of the results that follow. Let us denote the induced FABLE-posterior measure, the true data-generating measure, and the expectation under the true data-generating measure by $\widetilde{{\Pi}}$, $P_{0}$, and $E_{0}$, respectively.
@@
Suppose Assumptions 3.13.4 hold. Then, there exists a finite constant $G_{1}>0$ such that
@@
The proof of Proposition 3.6 is available in Section E.1 of the Supplementary Material. First, in Theorem 3.7, we provide FABLE-posterior contraction rates when estimating $\Lambda_{0}\Lambda_{0}^{\top},\Sigma_{0},$ and $\Psi_{0}$ using $\widetilde{L}=\widetilde{\Lambda}\widetilde{\Lambda}^{\top},\widetilde{\Sigma}$, and $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}$, respectively. Later, using Theorems 3.8 and 3.9, we justify the uncertainty quantification of the entrywise elements of $\Psi_{0}$ with the FABLE-posterior credible intervals.
3.2 FABLE-posterior Contraction Rates
@@
Suppose Assumptions 3.13.4 hold. Then, as $n\to\infty$, there exist finite constants $D_{1},D_{2},D>0$ such that
@@
    $E_{0}\left[\widetilde{\Pi}\left\{\dfrac{\|\widetilde{L}-\Lambda_{0}\Lambda_{0}^{\top}\|}{\|\Lambda_{0}\Lambda_{0}^{\top}\|}>D_{1}\left(\sqrt{\dfrac{\log n}{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
@@
    $E_{0}\left[\widetilde{\Pi}\left\{\|\widetilde{\Sigma}-\Sigma_{0}\|>D_{2}\left(\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}\right)\right\}\right]\to 0.$
@@
    $E_{0}\left[\widetilde{\Pi}\left\{\dfrac{\|\widetilde{\Psi}-\Psi_{0}\|}{\|\Psi_{0}\|}>D\left(\sqrt{\dfrac{\log n}{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
Theorem 3.7 shows the FABLE-posterior concentration of the relative errors around 0 when estimating $\Lambda_{0}\Lambda_{0}^{\top},\Sigma_{0},$ and $\Psi_{0}$. The result in part (b) is also a statement on relative error when estimating $\Sigma_{0}$, as $0<c_{2}\leq\|\Sigma_{0}\|=\max_{1\leq j\leq p_{n}}\,\sigma_{0j}^{2}=O(1)$ from Assumption 3.3. From part (c), the relative error in estimating the true covariance matrix converges to 0 at the rate of $n^{-1/2}+p_{n}^{-1/2}$, up to logarithmic factors. This showcases the blessing of dimensionality, as concentration of the FABLE-posterior is obtained if and only if both the number of samples $n$ and the number of dimensions $p_{n}$ increase. The proof of Theorem 3.7 is available in Section B of the Supplementary Material.